OpenStudy (anonymous):

I have a question on how my calculator handles logs, We all now that a log is kinda like the inverse of an exponent so to speak. and usually its written like 8=Log_2-3 but what is the calculator doing when type log 23 and thats it. and it spits out a decimal 1.36 what is that number . We didnt specify a base or what we want it to equal Any ideas

OpenStudy (anonymous):

i got 1.36 when i type in log23 as well

OpenStudy (anonymous):

your calculator has two logs, log base 10 and log base e if you want to find a log to a different base, you have to use the change of base formula

OpenStudy (anonymous):

so if you want \[\log_{10}(50)\] then you can type that in directly and get the answer in one step

OpenStudy (anonymous):

@satellite yea i get that but when you type just (LOG 23) what is it doing? its not log base 10 cuz that is not correct and its not LN cuz that is a seperate button

OpenStudy (anonymous):

but if you want \[\log_2(50)\] you have to type in \[\frac{\log(50)}{\log(2)}\] do find it

OpenStudy (anonymous):

\[\log(23)=1.3617...\] because \[10^{1.3617...}=23\]

OpenStudy (anonymous):

o so the "LOG" is automatically base 10 then

OpenStudy (anonymous):

but if you want \[\log_2(23)\] i.e. if you want to solve \[2^x=23\] for x you have to enter \[x=\frac{\log(23)}{\log(2)}\]

OpenStudy (anonymous):

yes "log" means \[\log_{10}\]

OpenStudy (anonymous):

similarly \[\ln\] means \[\log_e\]

OpenStudy (anonymous):

ok that is what i was trying to figure out, Thanks!

OpenStudy (anonymous):

so for example \[\log(1000)=3\] because \[10^3=1000\]

OpenStudy (anonymous):

yw

OpenStudy (zarkon):

Log always means natural long to me unless it specifies log base 10

OpenStudy (anonymous):

@zarkon log is log, but calculator ( pre ti 89) has log as a button, and that means \[\log_{10}\]

myininaya (myininaya):

natural long lol log always mean log base 10 to me and Ln means natural log

OpenStudy (zarkon):

lol...oops

OpenStudy (anonymous):

all logs are the same, just as all exponentials are the same. doesn't really matter what base is

OpenStudy (zarkon):

that is why we should always use base e :)

OpenStudy (anonymous):

that is why it never ceases to annoy me that math teachers want you to use the info given to write something as \[A_0e^{kt}\] when \[A_0 r^{\frac{t}{k}}\] would do just as well.

OpenStudy (anonymous):

care loses 15% of its value each year. why in gods green earth am i supposed to write the model for the cars value as \[A_0e^{kt}\] when i can use \[A_0(.85)^t\]?

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

for real. just because someone says i am supposed to use base e? and then due to rounding error i mess up all further calculations? makes no sense at all.